OpenStudy (anonymous):

giving medals for answers! Mary and Jane are both at the gas station at Hill and Elm. From the station, Mary drives 12 miles east and Jane drives 16 miles south. Find the distance between Mary and Jane.

OpenStudy (anonymous):

NVM!

OpenStudy (anonymous):

huh

OpenStudy (anonymous):



OpenStudy (anonymous):

there you go. now you have to figure out how far are they which is the hypotnuse.

OpenStudy (anonymous):

30

OpenStudy (anonymous):

20

OpenStudy (anonymous):

so a^2+b^2 is C^2

OpenStudy (anonymous):

i mean

OpenStudy (anonymous):

guessing is not gonna solve it TRUST ME

OpenStudy (anonymous):

want to do it step by step with me?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k so if a=12 and b=16 it would be 12^2 + 16^2=c^2 C is how far they are from eachother.

OpenStudy (anonymous):

when we square them it will look like: 144+256=c^2

OpenStudy (anonymous):

and that is 400

OpenStudy (anonymous):

400

OpenStudy (anonymous):

400=c^2 but we are looking for c so what do we do?

OpenStudy (anonymous):

we isolate the variable

OpenStudy (anonymous):

so \[\sqrt{400}=\sqrt{c}\]

OpenStudy (anonymous):

it would 4 sq root 28

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

sorry 20

OpenStudy (anonymous):

no atm the equation is just √400=√c^2

OpenStudy (anonymous):

so yeah 20 is correct

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

np!